if a 25ml of a solution of 50% is used then how much amount of 30% of same solution must be used?

Answer by ramkikk66(644)


25 ml of solution is at 50%. That means that it contains 50% of the solute e.g.
let us assume it is alcohol.
I.e. in a 50% solution of 25 ml, volume of alcohol = 25*50% = 25 * 0.5 = 12.5 ml
Now the question, if you had the same volume of alcohol, i.e. 12.5 ml, what is
the total volume of the mixture to make it a 30% solution.
If the 30% solution mixture has a volume of x ml, and 12.5 ml of it is alcohol,
then it means that:
x * 30% = x * 0.3 = 12.5
x = 12.5/0.3 = 41.67 ml approximately.
So to get the same amount of alcohol, you have to use 41.67 ml of a 30% solution.
Hope you got it :)